4+1+6x3=x which do you work this problem from lelf to right or right to lelf? when i went to school we work lelf to right , has this changed?
Found 2 solutions by richard1234, astevn816:
Answer by richard1234(7193)

Not sure what the x on the RHS of the equation is supposed to me...

The correct way is to use the order of operations, which states that operations within parentheses and exponents take precedence over multiplication and division, which take precedence over addition and subtraction (known as PEMDAS). You would do 6 x 3 = 18 first, and then add left-to-right: 4+1+18 = 23.

Answer by astevn816(5)

I guess the X is supposed to be a times symbol, anyway the order of operations call for BEMDAS. Since this puzzle is not in brackets you would do the multiplication first then the addition. 6 Multiplied by 3 = 18.
18 + 4 = 22
22 + 1 =23
the answer to this puzzle will be 23